# Precise Robotic Weed Spot-Spraying for Reduced Herbicide Usage and Improved Environmental Outcomes - A Real-World Case Study

### v1

@@ 6 Nutgrass Plant sugarcane 530,004 90,858
Table 1: Summary of images collected and annotated for each of the six spray trials.
Following data collection, a group of annotators manually annotated the presence of target weeds in each training dataset. Annotation was performed using the Computer Vision Annotation Tool (CVAT) cvat2020 . This allowed for rapid annotation of approximately 2,000 tile images per hour. Annotators labelled the presence of target weeds in each tile using a binary approach for spray trials 1, 2, 3 and 6. However, the training dataset for spray trials 4 and 5 required a multilabel classification approach due to the presence of two target weeds (grass and broadleaf) in the dataset. Fig. 6 provides sample images from the site-specific training datasets that were collected and annotated for each of the trials.
@@
3.3 DL training and validation for weed classification
@@
Following the labelling process, the training datasets for each trial were randomly split into 80% and 20% subsets of training and validation, where 80% of the labels were set aside for training and 20% of the labels were set aside for model validation. The class distribution between training and validation subsets was stratified such that the ratio of the weed class(es) to the negative class was consistent across both subsets.
@@
To allow for benchmarking of accuracy metrics for researchers, machine learning datasets are usually split three ways with training and validation subsets, and a test subset that is heldout to test the model. We have foregone this data splitting approach here in order to train models on the most data possible. Furthermore, the test set for this work is the real-time in situ data when evaluating spraying performance in the field.
Figure 7: Training and validation accuracy versus epoch (left axis) and training and validation loss versus epoch (right axis) during the training process for (a) spray trials 1 and 2, (b) spray trial 3, (c) spray trials 4 and 5 and (d) spray trial 6.
The TensorFlow machine learning backend was used together with the Python-based high level API Keras, to train a MobileNetV2 architecture following the training methodology of Olsen2019 ; Olsen2020 ; Calvert2021 . On average, the training process took $4-5$ hours on an NVIDIA GTX 1080Ti Graphical Processing Unit (GPU). The result of the training process is illustrated with time series plots of training and validation metrics versus successive epochs in Fig. 7.
Early stopping was used to halt training when the validation loss failed to decrease after 16 successive epochs. For each trial, the MobileNetV2 model from the epoch with the lowest validation loss was selected for field implementation. Fig. 8 presents the confusion matrices for each trial dataset which captures how well each model classifies its target weed(s).
Figure 8: The confusion matrices evaluating the performance of the MobileNetV2 weed detection models from (a) spray trials 1 and 2, (b) spray trial 3, (c) spray trials 4 and 5 and (d) spray trial 6.
@@
The light-weight MobileNetV2 architecture was chosen, similar to other works Olsen2019 ; Olsen2020 ; Calvert2021 , because it offers the high-speed inference required to achieve real-time spot spraying while travelling at up to 8 km per hour. When running on an NVIDIA Jetson Nano embedded device, the MobileNetV2 architecture performs inference at $21.9$ ms per image or $45.7$ frames per second.
@@ 3.4 Weed knockdown efficacy and herbicide usage analyses
To quantify the efficacy of each treatment on weeds, UAVs were deployed to capture high-resolution imagery from random locations along the trial paddock across all treatments before and after spraying. “Before maps” were collected on the day of the trial before applying the treatments. “After maps” were collected between 6-14 days after the treatment depending on the time it took visual symptoms to manifest on the targeted weeds after herbicide application. Fig. 10 presents the collected drone maps from each target trial site as part of the weed knockdown efficacy analysis. OpenDroneMap odm2020 was used to generate orthomosaics from the raw drone imagery and QGIS qgis2022 was used to georeference the before and after orthomosaics and extract comparative $1.5\times 1.5$ metre -- $1.6\times 1.6$ metre images from the maps. A team of annotators were then able to manually annotate each image counting the number of target weeds sprayed and target weeds missed. The weed knockdown hit rate for each treatment was calculated as
@@ 4 Results
$*{}^{*}$ Nutgrass / Ratoon sugarcane Sempra - - 200 177 - 11
@@ 6 Nutgrass / Plant sugarcane Sempra 100 96 207 73 96 65
Table 3: Summary of the spot spraying knockdown efficacy (Eq. 2) and herbicide usage reduction compared to blanket spraying results for the trials. $*{}^{*}$ote that no knockdown comparison was possible for all spray trial 1 plots due to technical error during the follow up UAV flights.
@@
$*{}^{*}$ Nutgrass / Ratoon sugarcane Sempra / Halosulfuron - - - - - -
$*{}^{*}$ Nutgrass / Ratoon sugarcane Sempra / Halosulfuron - - - - - -
${}^{\dagger}$ Nutgrass / Ratoon sugarcane Krismat / Ametryn 111.5 15 54.25 5.94 51% 60%
${}^{\dagger}$ Nutgrass / Ratoon sugarcane Krismat / Trifloxysulfuron 3.56 0.48 2.13 0.23 40% 51%
@@
Table 4: Summary of the water quality results for the trials. $*{}^{*}$ater quality analyses could not be completed for spray trials 1 and 2 due to more than 90mm of unforecasted rain in the days after spraying and before the planned date for water quality sampling. ${}^{\dagger}$pray trial 3 used Kristmat which has two active chemical agents present for the water quality analysis.
@@ 4.2 Herbicide usage reduction
Figure 15: A comparison of the herbicide usage results for blanket spraying and spot spraying across the trials (left axis). Along with the approximate weed density of the treated rows (right axis).
@@ 4.3 Water quality improvements
Water quality sampling and analysis was performed following the aforementioned methodology for trials 3, 4, 5, and 6. Table 4 documents the results and Figs. 16 and 17 compare the mean concentrations and loads of herbicide found in runoff between spot spraying and blanket spraying.
Figs. 16 and 17 show that herbicide concentrations and loads are proportional to the amount of herbicide sprayed across all trials. They also show that on average across all trials, spot spraying reduced the mean concentrations and loads of herbicides in runoff by 39% and 54% of the blanket spray, respectively. By limiting the amount of herbicide sprayed to areas with weed coverage and not spraying bare ground, spot spraying technology can reduce the environmental footprint of herbicide control.
@@ 5 Discussions and Future Works
In this study, we report a comprehensive set of field experiments testing robotic weed spot spraying when targeting different weeds and using different chemical herbicides. The results are very promising in terms of maintaining a high weed knockdown efficacy whilst significantly reducing the amount of herbicide applied and consequently the amount of herbicide detected in water run-off induced by irrigation. Therefore, robotic spot spraying has substantial potential to reduce herbicide use.
The scope of this work was limited to sugarcane paddocks with common weeds in the Burdekin region of Queensland, Australia. Therefore, the datasets and detection models collected and developed herein may not have utility in other sugarcane farming regions. Nevertheless, we expect that the general conclusions regarding the effectiveness of the technology would apply to other regions.
@@ 6 Conclusions
In conclusion, this work demonstrates the efficacy of a custom-designed robotic spot spraying tool, AutoWeed, in reducing herbicide usage on sugarcane farms while maintaining effective weed control.
Our field trials, spanning 25 hectares of sugarcane farms, show that AutoWeed spot spraying is 97% as effective at controlling weeds compared to the industry-standard broadcast spraying. Importantly, it also offers a substantial reduction in herbicide usage, with an average reduction of 35%, directly correlated with weed density. For areas with lower weed pressure, the reduction in herbicide usage can be as high as 65%.
Our analysis of irrigation-induced runoff also revealed that spot spraying reduced both the mean concentration and mean load of herbicides in runoff by 39% and 54%, respectively, compared to broadcast spraying. These findings underscore the potential of precise robotic spot spraying to benefit both farmers and the environment, by cutting input costs and providing sustained water quality improvements.

### v2

@@ 6 Nutgrass Plant sugarcane 530,004 90,858
Table 1: Summary of images collected and annotated for each of the six spray trials. Grass weeds included summer grass and crowsfoot; broadleaf weeds included vines, giant pigweed, and sesbania pea.
3.2.2 Dataset annotation for DL training
Following data collection, a group of annotators manually annotated the presence of target weeds in each weed training dataset in a binary fashion, i.e. labelled images with weed present as weed, and the others as non-target. A tiled classification approach was chosen whereby collected images were split into 2 $\times$ 2 tiles for annotation (Fig. 4a) and 1 $\times$ 2 tiles for inference during the use of the system for spot-spraying (Fig. 4b). This allowed rapid collection and annotation of high tile counts during dataset collection, and high-speed performance with a batch size of two during inference. It is worth noting that, the exact spatial position of the weeds was not recorded in annotations or used in our spray application. Instead, if an image tile contained a weed, its corresponding spray nozzle was activated, as shown in Fig. 4b. Because the design of the system is made for crops with rows that travel in a straight line direction, this system will accurately identify and spray weeds by only needing to tune the spray on time and spray duration based on the vehicle’s speed of travel.
Annotation was performed using the Computer Vision Annotation Tool (CVAT). This allowed for rapid annotation of approximately 2,000 tile images per hour, and an overall annotation time between 4.5 hours for trial 3 and 45 hours for trial 6. Annotators labelled the presence of target weeds in each tile using a binary approach for spray trials 1, 2, 3 and 6. However, the training dataset for spray trials 4 and 5 required a multi-label classification approach due to the presence of two target weeds (grass and broadleaf) in the dataset. Fig. 5 provides sample images from the site-specific training datasets that were collected and annotated for each of the trials.
@@ 3.2.2 Dataset annotation for DL training
3.2.3 DL training and validation for weed classification
@@
Following the labelling process, the training datasets for each trial were randomly split into 80% and 20% subsets of training and validation, where 80% of the labels were set aside for training and 20% of the labels were set aside for model validation. The class distribution between training and validation subsets was stratified such that the ratio of the weed class(es) to the negative class was (i.e. without weeds) consistent across both subsets.
@@
To allow for benchmarking of accuracy metrics for researchers, machine learning datasets are usually split three ways with training and validation subsets, and a test subset that is heldout to test the model. We have foregone this data splitting approach here, to train models on the most data possible. Furthermore, the test set for this work is the real-time in situ data when evaluating spot-spraying performance in the field.
The TensorFlow machine learning backend was used together with the Python-based high-level API Keras, to train a MobileNetV2 architecture following the training methodology of [6, 3]. MobileNetV2 is a deep learning model designed for efficient image classification tasks, particularly on mobile and embedded devices. It employs an architecture that balances high accuracy with low computational cost, using depthwise separable convolutions and inverted residuals. The loss function commonly used for training MobileNetV2 is categorical cross-entropy, which measures the difference between the predicted probability distribution and the true distribution of the class labels. On average, the training process took $4-5$ hours on an NVIDIA GTX 1080Ti Graphical Processing Unit (GPU). Early stopping was used to halt training when the validation loss failed to decrease after 16 successive epochs.
Fig. 6 presents the confusion matrices for each collected dataset, which captures how well each trained model classifies its target weed(s). This data shows lab inference using the MobileNetV2 model from the epoch with the lowest validation loss, which is also used for field implementation. We utilised 20% of the annotated data to test model performance shown in Fig. 6. The data volume used for the various trials can be calculated from the information provided in Table 1. This volume ranges from 1727 images (20% of 8638) to test the model developed for detecting nutgrass in trial 3, to 18172 images for testing the nutgrass detection model developed for trial 6.
Figure 6: The confusion matrices evaluating the performance of the MobileNetV2 weed detection models in lab inference on 20% of the labelled validation data. Negative (in a-b, and d) and Background (in c) refer to instances where there is no weed present in the image. The model may incorrectly identify an image with no weed as containing weed (false positive), or miss a weed (false negative). Here, (a) shows results from spray trials 1 and 2, (b) from spray trial 3, (c) from spray trials 4 and 5, and (d) from spray trial 6.
3.2.4 In-field spot-spraying trials
For each trial shown in Table 1, the MobileNetV2 model from the training process with the epoch with the lowest validation loss was selected for field implementation.
@@ 3.2.4 In-field spot-spraying trials
The light-weight MobileNetV2 architecture was chosen, similar to other works [6, 3], because it offers the high-speed inference required to achieve real-time spot spraying while travelling at up to 8 km per hour. When running on an NVIDIA Jetson Nano embedded device, the MobileNetV2 architecture performs inference at $21.9$ ms per image or $45.7$ frames per second.
Our in-field spot-spraying trials demonstrated that a model with high real-world inference accuracy is suitable for spot-spraying target weeds in the field for two main reasons. First, most field variations, such as ambient light, mechanical vibration, and background complexity, are captured in the large training dataset collected for each trial (see Table 1). Second, since the deep learning model has multiple opportunities to see the same weed due to the high frame processing rate and capturing images from the same view multiple times, the weed hit rate in the field is even higher than the per-image model accuracy. This is evidenced by the data presented in the next Section, which shows a high weed knockdown hit rate for spot-spraying, consistent with the accuracy of the deep learning models trained on the dataset.
There are also several sources of time delay from the moment that an image is captured to the time that the chemical hits the target weed. These include image acquisition time (i.e. the time it takes for the image to be captured and made available for processing), pre-processing time (i.e. the time it takes to prepare the image format for inference on the embedded GPU device), inference time (i.e. the time to perform inference using the embedded GPU target device and return a result indicating the presence of a weed or not) and solenoid response time (i.e. the time to send a command to engage a solenoid and have the solenoid electrically engaged).
For this work, we have measured the average time for each of these delays as shown in Table 2. The total average response time from image capture to weed spray is 58.16 ms, using the MobileNetV2 architecture. While travelling at 8 km/hr, the spray vehicle would have only moved approximately 129.2mm over this response time. To ensure adequate coverage of the target weed, we activate the sprayer as quickly as possible to ensure that the sprayer is fully on before the centre of the sprayer passes over the identified weed target. We also instituted a spray duration, that is based on the vehicle speed, so that the sprayer remains on until the sprayer has passed over the target weed to give it full coverage of herbicide. Typically, a spray duration of 0.45 seconds was used for a travel speed of 8 km/hr. This ensured spray sections of approximately 1 metre for identified targets.
Time Measurement Average (ms) Std. Dev. (ms)
Image acquisition time 5.85 0.75
Image pre-processing time 8.88 0.05
Inference time 21.90 5.53
Solenoid response time 21.53 1.70
Total execution time 58.16 5.83
Table 2: Average time measurements for the spot spraying workflow from image acquisition to solenoid activation showing the average and standard deviation measurements in milliseconds.
@@ 3.3 Weed knockdown efficacy and herbicide usage analyses
To quantify the efficacy of each treatment on weeds, UAVs were deployed to capture high-resolution imagery from random locations along the trial paddock across all treatments before and after spraying. “Before maps” were collected on the day of the trial before applying the treatments. “After maps” were collected between 6-14 days after the treatment depending on the time it took visual symptoms to manifest on the targeted weeds after herbicide application. Fig. 8 presents the collected drone maps from each target trial site as part of the weed knockdown efficacy analysis. OpenDroneMap [28] was used to generate orthomosaics from the raw drone imagery and QGIS [29] was used to georeference the before and after orthomosaics and extract comparative $1.5\times 1.5$ metre -- $1.6\times 1.6$ metre images from the maps. A team of annotators were then able to manually annotate each image counting the number of target weeds sprayed and target weeds missed. The weed knockdown hit rate for each treatment was calculated as
@@ 4 Results
1 Nutgrass / Ratoon sugarcane Sempra - - 200 177 - 11
@@ 6 Nutgrass / Plant sugarcane Sempra 100 96 207 73 96 65
Table 4: Summary of the spot spraying knockdown efficacy (Eq. 2) and herbicide usage reduction compared to blanket spraying results for the trials. Note that no knockdown comparison was possible for all spray trial 1 plots due to technical error during the follow up UAV flights.
@@
1 Nutgrass / Ratoon sugarcane Sempra / Halosulfuron - - - - - -
2 Nutgrass / Ratoon sugarcane Sempra / Halosulfuron - - - - - -
3 Nutgrass / Ratoon sugarcane Krismat / Ametryn 111.5 15 54.25 5.94 51% 60%
3 Nutgrass / Ratoon sugarcane Krismat / Trifloxysulfuron 3.56 0.48 2.13 0.23 40% 51%
@@ 2 Nutgrass / Ratoon sugarcane Sempra / Halosulfuron - - - - - -
Table 5: Summary of the water quality results for the trials. Water quality analyses could not be completed for spray trials 1 and 2 due to more than 90mm of unforecasted rain in the days after spraying and before the planned date for water quality sampling. Spray trial 3 used Kristmat which has two active chemical agents present for the water quality analysis.
@@ 4.2 Herbicide usage reduction
Figure 13: A comparison of herbicide usage (left axis) for blanket spraying and spot spraying across the trials. Weed density (%) is measured as the number of images with a weed detected in them divided by the total number of images in each trial area.
4.3 Statistical analysis of spray trial data
This section presents the statistical analysis of the spray trial data, comparing spot and blanket herbicide application methods. The analysis encompasses herbicide usage rates and the correlation between weed density and spot spray application rate.
4.3.1 Herbicide usage analysis
We calculated the mean and standard deviation of herbicide used in L/Ha for both spot and blanket spray methods. This analysis provides insights into the overall herbicide usage and its variability across different plots. As shown in Table 6, the spot-spraying method shows much higher variability (higher Standard Deviation), which is attributed to the highly variable weed density (see Fig. 13) in the various plots in the 6 trials.
To determine if there is a significant difference in herbicide usage data between spot and blanket spray methods, we conducted Welch’s t-test, which is appropriate for samples with unequal variances. The p-value shows a statistically significant difference between the herbicide usage using two spraying methods.
4.3.2 Correlation analysis
Fig. 13 illustrates a clear positive correlation between weed density and herbicide application rate. To quantify this relationship, we calculated the Pearson correlation coefficient, yielding a value of $r=0.9824$, as shown in Table 6, indicating a very strong correlation between the two variables. This also explains the high standard deviation of herbicide usage in spot spraying trials, as weed density varies randomly across different trial paddocks.
Table 6: Statistical Analysis of Spray Trial Data
Metric Spot Spray Blanket Spray Test Statistic p-value
Mean SD Mean SD
Herbicide Usage (L/Ha) 132 50 204 9 $t=-4.5754^{a}$ 0.0007
Weed Knockdown (%) 94.67 4.44 98.33 1.37 - -
Correlation: Weed Density vs. Herbicide Used $r=0.9824^{b}$
a Paired t-test to compare Herbicide Usage rate between Spot and Blanket methods
b Pearson correlation coefficient
@@ 4.4 Water quality improvements
Water quality sampling and analysis was performed following the aforementioned methodology for trials 3, 4, 5, and 6. Table 5 documents the results and Figs. 14 and 15 compare the mean concentrations and loads of herbicide found in runoff between spot spraying and blanket spraying.
Figs. 14 and 15 show that herbicide concentrations and loads are proportional to the amount of herbicide sprayed across all trials. They also show that on average across all trials, spot spraying reduced the mean concentrations and loads of herbicides in runoff by 39% and 54% of the blanket spray, respectively. By limiting the amount of herbicide sprayed to areas with weed coverage and not spraying bare ground, spot spraying technology can reduce the environmental footprint of herbicide control.
@@
Paired t-tests on trial averages indicate a trend of reduced knockdown (Fig. 12), concentration (Fig. 14), and loads (Fig. 15) with spot spraying, aligning with expectations. However, the small sample sizes and small number of within-trial replicates limit statistical power.
@@ 5 Discussions and Future Works
In this study, we report a comprehensive set of field experiments testing robotic weed spot spraying when targeting different weeds and using different chemical herbicides. The results are very promising in terms of maintaining a high weed knockdown efficacy whilst significantly reducing the amount of herbicide applied and consequently, the amount of herbicide detected in water run-off induced by irrigation. Therefore, robotic spot spraying has substantial potential to reduce herbicide use.
The technology presented in this project was specifically developed for managing sugarcane weeds. However, crop rotation in sugarcane paddocks is a fundamental practice in sustainable agriculture, contributing to long-term soil productivity, environmental health, and economic viability for farmers. We demonstrated that our technology is effective even when different rotation crops are cultivated. This adaptability is a strength of our spot-spraying technology that can seamlessly adjust to other weeds and crops using the same detection and spraying hardware, as well as the same detection framework. Our DL detection method remains consistent across the five different paddocks (with varying growth and natural environments), three different weed categories (with diverse morphologies, colours, and textures), and the two different crops (with distinct morphologies, colours, and textures). Additionally, our trial 6 was conducted in a plant cane rather than a ratoon cane, adding extra variability to our experiments. This approach is aligned with the well-known site-specific approach, one needs to retrain a previous model or develop a new model using data from a new paddock, which could have a previously addressed or a new weed/crop.
The scope of this work was limited to sugarcane (and a rotation crop in sugarcane) paddocks with common weeds in the Burdekin region of Queensland, Australia. Therefore, the datasets and detection models collected and developed herein may not have utility in other sugarcane farming regions, and a site-specific approach should be adopted for new weed/crop scenarios. Nevertheless, we expect that the general conclusions regarding the effectiveness of the technology would apply to other regions.
Another limitation of our work is the lack of yield data for our trials. This was primarily because yield analysis was beyond the scope of the current study, as we did not have the resources to obtain and analyse yield data for the farm blocks where we conducted our spraying trials. However, it is known that registered herbicides can negatively impact crop yield. Sugar Research Australia routinely screens a range of herbicides for their impact on cane varieties, with results updated yearly and available in the Sugar Research Australia variety guides [30]. When weeds are controlled and not competing with the crop, reduced herbicide usage may improve yield. Since we have demonstrated that spot-spraying with AutoWeed can reduce herbicide use while maintaining weed control, we expect that AutoWeed may positively impact overall yield. This needs to be fully verified in future studies.
@@
Another limitation of our study is the slight unintentional variations in our experiments, which occurred due to the nature of trial work involving several uncontrollable factors. Ideally, we would have conducted all trials on sugarcane. However, only four out of six trials were performed on sugarcane, with the remaining two conducted on a rotation crop on a sugarcane farm to utilise all available opportunities. Additionally, we had to use two different spraying systems due to the unavailability of the initial farming machine. Nonetheless, this did not impact our results as the detection and spraying control electronics remained consistent.
@@ 6 Conclusions
This study highlights how advanced robotics combined with computer vision and deep learning can revolutionise weed control in agriculture. The field evaluation of AutoWeed shows that precision application not only reduces herbicide usage on sugarcane farms but also promotes a more sustainable approach by minimising chemical use.
By targeting only the necessary areas, AutoWeed shifts the focus from blanket application to a smarter method, cutting input costs and reducing the chemical burden on ecosystems. The resultant improvements in water quality further underscore the environmental benefits of this technology.
In summary, AutoWeed represents a significant step forward in precision agriculture, setting the stage for future innovations that marry technology with sustainability to safeguard both farm productivity and the environment.